Dorcan Corporation manufactures and sells T-shirts imprinted with college names and slogans. Last year, the shirts sold for $8.00 each, and the variable cost to manufacture them was $3.75 per unit. The company needed to sell 20,000 shirts to break-even. The after tax net income last year was $6,180. Donnelly's expectations for the coming year include the following: (CMA adapted) The sales price of the T-shirts will be $10. Variable cost to manufacture will increase by one-third. Fixed costs will increase by 10%. The income tax rate of 40% will be unchanged. Based on a $10 selling price per unit and if Dorcan Corporation wishes to earn $81,987 in after tax net income for the coming year, the company's sales volume in dollars must be: Multiple Choice $45,389. $46,829. $44,429. $460,290.
Answer #1

Fixed cost = Break even*Contribution margin per unit = 20000*4.25 = 85000

Present variable cost = 3.75+(3.75/3) = 5

Present fixed cost = 85000*1.1 = 93500

Income before tax = 81987*100/60 = 136645

Required sales volume = (136645+93500)/5 = 46029

Required sales = 46029*10 = 460290

So answer is d) $460290